OpenStudy (patrickweewoos):

Simplify the following expression. (3^3)^-1

OpenStudy (patrickweewoos):

@iGreen

OpenStudy (gorv):

3^3=3*3*3=???

OpenStudy (igreen):

\((3^3)^{-1}\) Multiply the exponents. What's 3 * -1? @PatrickWeeWoos

OpenStudy (patrickweewoos):

-3

OpenStudy (anonymous):

3^-3 = ?

OpenStudy (anonymous):

Wait, that came out wrong...

OpenStudy (igreen):

Yep, that gives us \(3^{-3}\). Now what's 3^-3? Type it in your calculator.

OpenStudy (patrickweewoos):

9

OpenStudy (anonymous):

...

OpenStudy (anonymous):

No...

OpenStudy (igreen):

You can just put it in google.. No that's 3 * 3.

OpenStudy (igreen):

Put '3^-3' in google..it'll give you the answer..

OpenStudy (patrickweewoos):

1/27

OpenStudy (igreen):

Yes!

OpenStudy (igreen):

That's the fraction form of 3^-3.